# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
respectively. In (SI.5.2), $A_{\mathrm{tr}}$ and $E_{\mathrm{tr}}$ are the matrices defined in (SI.4.2) and (SI.4.3).
@@
Let $\Gamma^{\ast}$ be the set of parameters learned from the ridge regression problem in (20). Let $A_{\mathrm{test}}\in\mathbb{R}^{d\times(d+1)}$ and $B_{\mathrm{test}}\in\mathbb{R}^{(d+1)\times(d+1)}$ be two matrices constructed as in (SI.5.1) or (SI.5.2). We have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})\simeq\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{test}}^{\top}\right),$ (SI.5.3)
@@
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top})\simeq\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{T})-\frac{c_{e}}{d}\operatorname{tr}\left(B_{\mathrm{test}}\left[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}\right]\right).$ (SI.5.4)
@@
$\Gamma^{\ast}_{\mathrm{eq}}$ $\coloneqq\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}(E_{\mathrm{tr}}+\xi I)^{-1},$ (SI.5.5)
@@
$\xi\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\frac{\tau\lambda}{\xi}=1-\tau,$ (SI.5.6)
and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (SI.8.3). For convenience, this can be recast so that $\xi$ is the positive double root of
@@
Moreover, the scalar $c_{e}$ in (SI.5.4) is defined as
$c_{e}=\frac{\rho+\nu-\nu^{2}\mathcal{M}_{\kappa}(\nu)-\xi\left[1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right]}{1-{2\xi}\mathcal{M}_{\kappa}(\nu)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)-\tau},$ (SI.5.7)
@@
$\nu\coloneqq\frac{1+\rho}{\alpha}+\xi.$ (SI.5.8)
To derive the asymptotic characterizations (SI.5.3) and (SI.5.4) in Result 4, we first use block-matrix inversion to rewrite $\mathcal{G}_{e}(\pi)$ in (SI.4.8) as
${\mathcal{G}}_{e}(\pi)=\begin{bmatrix}c^{\ast}(\pi)&-c^{\ast}(\pi)\,(q^{\ast}(\pi))^{\top}\\ -c^{\ast}(\pi)\,q^{\ast}(\pi)&I\otimes F_{E}(\chi_{\pi})+c^{\ast}(\pi)q^{\ast}(\pi)(q^{\ast}(\pi))^{\top}\end{bmatrix},$ (SI.5.9)
where $F_{E}(\cdot)$ is the matrix-valued function defined in (SI.4.36), i.e.,
$F_{E}(\chi_{\pi})=\Big{(}\frac{\tau}{1+\chi_{\pi}}E_{\mathrm{tr}}+\pi B_{\mathrm{test}}+\lambda\tau I_{d+1}\Big{)}^{-1}.$ (SI.5.10)
@@
$q^{\ast}(\pi)=\frac{\tau}{(1+\chi_{\pi})\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\chi_{\pi})\right),$ (SI.5.11)
@@
$\frac{1}{c^{\ast}(\pi)}=\frac{\tau(1+\rho)}{1+\chi_{\pi}}+\lambda\tau-\frac{\tau^{2}}{(1+\chi_{\pi})^{2}d}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\chi_{\pi})\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}^{\top}\right).$ (SI.5.12)
Observe that there is a one-to-one correspondence between the terms in (SI.5.9) and those in (SI.3.8).
To derive the asymptotic characterization given in (SI.5.3), we note that
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})$ $\simeq\frac{-1}{c(0)\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{test}})^{T}\end{bmatrix}\mathcal{G}_{e}(0)e_{1}$ (SI.5.13)
$=\frac{c^{\ast}(0)}{c(0)}\cdot\frac{1}{d}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}A_{\mathrm{test}}^{\top}\right)$ (SI.5.14)
$\simeq\frac{1}{d}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}A_{\mathrm{test}}^{\top}\right).$ (SI.5.15)
In the above display, (SI.5.13) follows from (SI.3.12) and the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. The equality in (SI.5.14) is due to (SI.5.9) and (SI.5.11). To reach (SI.5.15), we note that $c(0)=e_{1}^{\top}G_{e}(0)e_{1}$ and $c^{\ast}(0)=e_{1}^{\top}\mathcal{G}_{e}(0)e_{1}$. Thus, $c(0)\simeq c^{\ast}(0)$ due to the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. In Appendix SI.8, we show that
$\lambda(1+\chi_{0})\simeq\xi,$ (SI.5.16)
where $\xi$ is the scalar defined in (SI.5.6). The asymptotic characterization given in (SI.5.3) then follows from (SI.5.15) and from the definition of $\Gamma^{\ast}_{\mathrm{eq}}$ given in (SI.5.5).
Next, we use (SI.3.14) to derive the asymptotic characterization of the quadratic term in (SI.5.4). Taking the derivative of (SI.5.12) gives us
@@
$\qquad-\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}\left(1+\rho-\frac{2}{d}\operatorname{tr}(A_{\mathrm{tr}}(E_{\mathrm{tr}}+\xi I)^{-1}A_{\mathrm{tr}}^{T})+\frac{1}{d}\operatorname{tr}(A_{\mathrm{tr}}(E_{\mathrm{tr}}+\xi I)^{-1}E_{\mathrm{tr}}(E_{\mathrm{tr}}+\xi I)^{-1}A_{\mathrm{tr}}^{T})\right)$ (SI.5.17)
$=\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})-\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}\left(1+\rho-\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{T})-\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})\right),$ (SI.5.18)
where $A_{\mathrm{tr}}$ is the matrix defined in (SI.4.2). In reaching the above expression, we have also used the estimate in (SI.5.16).
@@
$A_{\mathrm{tr}}=S\left(E_{\mathrm{tr}}+\xi I_{d+1}-\Big{(}\frac{1+\rho}{\alpha}+\xi\Big{)}I_{d+1}\right),$ (SI.5.19)
where $S$ is a $d\times(d+1)$ matrix obtained by removing the last row of $I_{d+1}$. Using this identity, we can rewrite the matrix $\Gamma^{\ast}_{\mathrm{eq}}$ in (SI.5.5) as
$\Gamma^{\ast}_{\mathrm{eq}}$ $=S\left(I-\Big{(}\frac{1+\rho}{\alpha}+\xi\Big{)}(E_{\mathrm{tr}}+\xi I)^{-1}\right)$ (SI.5.20)
$=\begin{bmatrix}I-\nu F_{R}(\nu)-a^{\ast}(1+\rho)^{2}\nu F_{R}(\nu)b_{\mathrm{tr}}b_{\mathrm{tr}}^{\top}F_{R}(\nu)&a^{\ast}(1+\rho)\nu F_{R}(\nu)b_{\mathrm{tr}}\end{bmatrix},$ (SI.5.21)
where $F_{R}(\cdot)$ is the function defined in (SI.8.1), and $\nu$ is the parameter given in (SI.5.8). The second equality (SI.5.21) is obtained from the explicit formula for $(E_{\mathrm{tr}}+\xi I)^{-1}$ in (SI.8.7).
From (SI.5.19) and (SI.5.20), it is straightforward to check that
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{T})=1-\nu+\nu^{2}\frac{1}{d}\operatorname{tr}(S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top}),$ (SI.5.22)
@@
$\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})=\xi\left[1-2\nu\frac{1}{d}\operatorname{tr}(S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top})+\nu^{2}\frac{1}{d}\operatorname{tr}(S(E_{\mathrm{tr}}+\xi I)^{-2}S^{\top}\right].$ (SI.5.23)
By using the asymptotic characterizations given in (SI.8.19) and (SI.8.20), we then have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{T})\simeq 1-\nu+\nu^{2}\mathcal{M}_{\kappa}(\nu),$ (SI.5.24)
@@
$\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})\simeq\xi\left[1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right].$ (SI.5.25)
Substituting (SI.5.24), (SI.5.25), and (SI.8.21) into (SI.5.18) yields
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c^{\ast}(\pi)}\right)\biggr{|}_{\pi=0}\simeq\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{T})-\frac{c_{e}}{d}\operatorname{tr}\left(B_{\mathrm{test}}\left[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}\right]\right),$ (SI.5.26)
where $c_{e}$ is the scalar defined in (SI.5.7). The asymptotic characterization of the quadratic term in (SI.5.4) then follows from (SI.3.14) and the claim that
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr{|}_{\pi=0}\simeq\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c^{\ast}(\pi)}\right)\biggr{|}_{\pi=0}.$ (SI.5.27)
SI.5.2 The Generalization Error of In-Context Learning

### v3

@@ Appendix D An Asymptotic Equivalent of the Extended Resolvent Matrix
Using (134), we can now write the equation (133) as
$\widehat{\chi}_{\mathrm{ave}}$ $\simeq\frac{1}{d}\operatorname{tr}\left(F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}\right)$ (138)
[…]
The second term on the right-hand side of (138) is negligible. Indeed,
[…]
$\qquad\qquad\qquad\leq\mathinner{\!\left\lVert F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}F_{E}(\widehat{\chi}_{\mathrm{ave}})\right\rVert}_{\mathsf{op}}(\,\mathinner{\!\left\lVert R_{\mathrm{tr}}\right\rVert}_{\mathsf{F}}^{2}+(1+\rho)^{2}\mathinner{\!\left\lVert b_{\mathrm{tr}}\right\rVert}^{2}).$
By construction, $\mathinner{\!\left\lVert F_{E}(\widehat{\chi}_{\mathrm{ave}})\right\rVert}_{\mathsf{op}}\leq(\lambda\tau)^{-1}$. Moreover, since the task vectors $\left\{w_{i}\right\}_{i\in[k]}$ are independent vectors sampled from $\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$, it is easy to verify that
$\mathinner{\!\left\lVert E_{\mathrm{tr}}\right\rVert}_{\mathsf{op}}=\mathcal{O}_{\prec}(1),\qquad\mathinner{\!\left\lVert R_{\mathrm{tr}}\right\rVert}_{\mathsf{F}}=\mathcal{O}_{\prec}(\sqrt{d})\qquad\text{and}\qquad\mathinner{\!\left\lVert b_{\mathrm{tr}}\right\rVert}_{2}=\mathcal{O}_{\prec}(1).$ (140)
Finally, since $\bar{c}$ is an element of $\widehat{\mathcal{G}}_{e}$, we must have $\mathinner{\!\left\lvert\bar{c}\right\rvert}\leq\mathinner{\!\left\lVert\widehat{\mathcal{G}}_{e}\right\rVert}_{\mathsf{op}}\leq(\tau\lambda)^{-1}$. Combining these estimates gives us
$\frac{\bar{c}\,\tau^{2}}{(1+\widehat{\chi}_{\mathrm{ave}})^{2}d^{3}}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}F_{E}(\widehat{\chi}_{\mathrm{ave}})\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}^{\top}\right)=\mathcal{O}_{\prec}(d^{-2}),$ (141)
and thus we can simplify (138) as
$\widehat{\chi}_{\mathrm{ave}}\simeq\frac{1}{d}\operatorname{tr}\left[\Big(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}E_{\mathrm{tr}}+\pi B+\lambda\tau I_{d}\Big)^{-1}E_{\mathrm{tr}}\right].$ (142)
Observe that (142) is a small perturbation of the self-consistent equation in (106). By the stability of the equation (106), we then have
$\widehat{\chi}_{\mathrm{ave}}\simeq\chi_{\pi},$ (143)
where $\chi_{\pi}$ is the unique positive solution to (106).
Recall the definitions of $\mathcal{G}_{e}(\pi)$ and $\widehat{\mathcal{G}}_{e}(\pi)$ in (130) and (107), respectively. By the standard resolvent identity,
$\widehat{\mathcal{G}}_{e}(\pi)-\mathcal{G}_{e}(\pi)$ (144)
By construction, $\mathinner{\!\left\lVert\widehat{\mathcal{G}}_{e}(\pi)\right\rVert}_{\mathsf{op}}\leq 1/(\tau\lambda)$ and $\mathinner{\!\left\lVert\mathcal{G}_{e}(\pi)\right\rVert}_{\mathsf{op}}\leq 1/(\tau\lambda)$. Moreover, $\mathinner{\!\left\lVert E_{\mathrm{tr}}\right\rVert}_{\mathsf{op}}\prec 1$ and
$\mathinner{\!\left\lVert\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)\right\rVert}\prec 1.$ (145)
It then follows from (143) and (144) that
$\mathinner{\!\left\lVert\widehat{\mathcal{G}}_{e}(\pi)-\mathcal{G}_{e}(\pi)\right\rVert}_{\mathsf{op}}\simeq 0.$ (146)
If $\widehat{\mathcal{G}}_{e}(\pi)$ satisfies the equivalent conditions (103), (104) and (105) (as claimed in our analysis above), then the estimate in (146) allows us to easily check that $\mathcal{G}_{e}(\pi)$ also satisfies (103), (104) and (105). Thus, we claim that $\mathcal{G}_{e}(\pi)$ is asymptotically equivalent to the extended resolvent matrix $G_{e}(\pi)$ in the sense of Definition 1.
Appendix E Asymptotic Limits of the Generalization Errors
E.1 Asymptotic Limits of the Linear and Quadratic Terms
From Corollary 1 and the discussions in Remark 2, characterizing the test error $e(\Gamma^{\ast})$ boils down to computing the linear term $\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}A_{\mathrm{test}}^{\top}\right)$ and the quadratic term $\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top}\right)$, where $A_{\mathrm{test}}$ and $B_{\mathrm{test}}$ are the matrices defined in (70) and (71), respectively.
We consider two different types of test data distributions $\mathcal{P}_{\mathrm{test}}$: ICL and IDG. [See the main text for details.] From (62) and (63), these two settings correspond to choosing
$\text{(ICL)}\mathrel{\mathop{\ordinarycolon}}\qquad A_{\mathrm{test}}=\begin{bmatrix}I_{d}&0\end{bmatrix}\quad\text{and}\quad B_{\mathrm{test}}=\begin{bmatrix}(\frac{1+\rho}{\alpha}+1)I_{d}&\\ &(1+\rho)^{2}\end{bmatrix}$. (147)
and
$\text{(IDG)}\mathrel{\mathop{\ordinarycolon}}\qquad A_{\mathrm{test}}=A_{\mathrm{tr}}\quad\text{and}\quad B_{\mathrm{test}}=E_{\mathrm{tr}},$ (148)
respectively. In (148), $A_{\mathrm{tr}}$ and $E_{\mathrm{tr}}$ are the matrices defined in (101) and (102).
@@ E.1 Asymptotic Limits of the Linear and Quadratic Terms
Let $\Gamma^{\ast}$ be the set of parameters learned from the ridge regression problem in (77). Let $A_{\mathrm{test}}\in\mathbb{R}^{d\times(d+1)}$ and $B_{\mathrm{test}}\in\mathbb{R}^{(d+1)\times(d+1)}$ be two matrices constructed as in (147) or (148). We have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})\simeq\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{test}}^{\top}\right),$ (149)
@@
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top})\simeq\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{T})-\frac{c_{e}}{d}\operatorname{tr}\left(B_{\mathrm{test}}\left[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}\right]\right).$ (150)
@@
$\Gamma^{\ast}_{\mathrm{eq}}$ $\coloneqq\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}(E_{\mathrm{tr}}+\xi I)^{-1},$ (151)
@@
$\xi\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi\right)-\frac{\tau\lambda}{\xi}=1-\tau,$ (152)
and $\mathcal{M}_{\kappa}(\cdot)$ is the function defined in (252). For convenience, this can be recast so that $\xi$ is the positive double root of
@@
Moreover, the scalar $c_{e}$ in (150) is defined as
$c_{e}=\frac{\rho+\nu-\nu^{2}\mathcal{M}_{\kappa}(\nu)-\xi\left[1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right]}{1-{2\xi}\mathcal{M}_{\kappa}(\nu)-\xi^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)-\tau},$ (153)
@@
$\nu\coloneqq\frac{1+\rho}{\alpha}+\xi.$ (154)
To derive the asymptotic characterizations (149) and (150) in Result 4, we first use block-matrix inversion to rewrite $\mathcal{G}_{e}(\pi)$ in (107) as
${\mathcal{G}}_{e}(\pi)=\begin{bmatrix}c^{\ast}(\pi)&-c^{\ast}(\pi)\,(q^{\ast}(\pi))^{\top}\\ -c^{\ast}(\pi)\,q^{\ast}(\pi)&I\otimes F_{E}(\chi_{\pi})+c^{\ast}(\pi)q^{\ast}(\pi)(q^{\ast}(\pi))^{\top}\end{bmatrix}$, (155)
where $F_{E}(\cdot)$ is the matrix-valued function defined in (135), i.e.,
$F_{E}(\chi_{\pi})=\Big(\frac{\tau}{1+\chi_{\pi}}E_{\mathrm{tr}}+\pi B_{\mathrm{test}}+\lambda\tau I_{d+1}\Big)^{-1}.$ (156)
@@
$q^{\ast}(\pi)=\frac{\tau}{(1+\chi_{\pi})\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\chi_{\pi})\right),$ (157)
@@
[…]
Observe that there is a one-to-one correspondence between the terms in (155) and those in (87).
To derive the asymptotic characterization given in (149), we note that
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}A_{\mathrm{test}}^{\top})$ $\simeq\frac{-1}{c(0)\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{test}})^{T}\end{bmatrix}\mathcal{G}_{e}(0)e_{1}$ (159)
$=\frac{c^{\ast}(0)}{c(0)}\cdot\frac{1}{d}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}A_{\mathrm{test}}^{\top}\right)$ (160)
$\simeq\frac{1}{d}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}A_{\mathrm{test}}^{\top}\right).$ (161)
In the above display, (159) follows from (91) and the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. The equality in (160) is due to (155) and (157). To reach (161), we note that $c(0)=e_{1}^{\top}G_{e}(0)e_{1}$ and $c^{\ast}(0)=e_{1}^{\top}\mathcal{G}_{e}(0)e_{1}$. Thus, $c(0)\simeq c^{\ast}(0)$ due to the asymptotic equivalence between $G_{e}(0)$ and $\mathcal{G}_{e}(0)$. In Appendix H, we show that
$\lambda(1+\chi_{0})\simeq\xi,$ (162)
where $\xi$ is the scalar defined in (152). The asymptotic characterization given in (149) then follows from (161) and from the definition of $\Gamma^{\ast}_{\mathrm{eq}}$ given in (151).
Next, we use (93) to derive the asymptotic characterization of the quadratic term in (150). Taking the derivative of (158) gives us
@@
$\qquad-\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}\left(1+\rho-\frac{2}{d}\operatorname{tr}(A_{\mathrm{tr}}(E_{\mathrm{tr}}+\xi I)^{-1}A_{\mathrm{tr}}^{T})+\frac{1}{d}\operatorname{tr}(A_{\mathrm{tr}}(E_{\mathrm{tr}}+\xi I)^{-1}E_{\mathrm{tr}}(E_{\mathrm{tr}}+\xi I)^{-1}A_{\mathrm{tr}}^{T})\right)$ (163)
$=\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})-\frac{\tau\chi^{\prime}_{0}}{(1+\chi_{0})^{2}}\left(1+\rho-\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{T})-\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})\right),$ (164)
where $A_{\mathrm{tr}}$ is the matrix defined in (101). In reaching the above expression, we have also used the estimate in (162).
@@
$A_{\mathrm{tr}}=S\left(E_{\mathrm{tr}}+\xi I_{d+1}-\Big(\frac{1+\rho}{\alpha}+\xi\Big)I_{d+1}\right),$ (165)
where $S$ is a $d\times(d+1)$ matrix obtained by removing the last row of $I_{d+1}$. Using this identity, we can rewrite the matrix $\Gamma^{\ast}_{\mathrm{eq}}$ in (151) as
$\Gamma^{\ast}_{\mathrm{eq}}$ $=S\left(I-\Big(\frac{1+\rho}{\alpha}+\xi\Big)(E_{\mathrm{tr}}+\xi I)^{-1}\right)$ (166)
$=\begin{bmatrix}I-\nu F_{R}(\nu)-a^{\ast}(1+\rho)^{2}\nu F_{R}(\nu)b_{\mathrm{tr}}b_{\mathrm{tr}}^{\top}F_{R}(\nu)&a^{\ast}(1+\rho)\nu F_{R}(\nu)b_{\mathrm{tr}}\end{bmatrix}$, (167)
where $F_{R}(\cdot)$ is the function defined in (250), and $\nu$ is the parameter given in (154). The second equality (167) is obtained from the explicit formula for $(E_{\mathrm{tr}}+\xi I)^{-1}$ in (256).
From (165) and (166), it is straightforward to check that
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{T})=1-\nu+\nu^{2}\frac{1}{d}\operatorname{tr}(S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top}),$ (168)
@@
$\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})=\xi\left[1-2\nu\frac{1}{d}\operatorname{tr}(S(E_{\mathrm{tr}}+\xi I)^{-1}S^{\top})+\nu^{2}\frac{1}{d}\operatorname{tr}(S(E_{\mathrm{tr}}+\xi I)^{-2}S^{\top}\right].$ (169)
By using the asymptotic characterizations given in (268) and (269), we then have
$\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}A_{\mathrm{tr}}^{T})\simeq 1-\nu+\nu^{2}\mathcal{M}_{\kappa}(\nu),$ (170)
@@
$\frac{\xi}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}(\Gamma^{\ast}_{\mathrm{eq}})^{\top})\simeq\xi\left[1-2\nu\mathcal{M}_{\kappa}(\nu)-\nu^{2}\mathcal{M}^{\prime}_{\kappa}(\nu)\right].$ (171)
Substituting (170), (171), and (270) into (164) yields
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c^{\ast}(\pi)}\right)\biggr|_{\pi=0}\simeq\frac{1}{d}\operatorname{tr}(\Gamma^{\ast}_{\mathrm{eq}}B_{\mathrm{test}}(\Gamma^{\ast}_{\mathrm{eq}})^{T})-\frac{c_{e}}{d}\operatorname{tr}\left(B_{\mathrm{test}}\left[(E_{\mathrm{tr}}+\xi I)^{-1}-\xi(E_{\mathrm{tr}}+\xi I)^{-2}\right]\right),$ (172)
where $c_{e}$ is the scalar defined in (153). The asymptotic characterization of the quadratic term in (150) then follows from (93) and the claim that
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c(\pi)}\right)\biggr|_{\pi=0}\simeq\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\pi}\left(\frac{1}{c^{\ast}(\pi)}\right)\biggr|_{\pi=0}.$ (173)
E.2 The Generalization Error of In-Context Learning